# OceanSplat: Object-aware Gaussian Splatting with Trinocular View Consistency for Underwater Scene Reconstruction

### v1

@@ Introduction
We evaluate OceanSplat with previous methods, including the official and Nerfstudio versions of SeaThru-NeRF (levy2023seathru), as well as 3DGS (kerbl20233d), SeaSplat (yang2024seasplat), and WaterSplatting (li2024watersplatting). All results were reproduced using the publicly released code. As shown in Table 1, OceanSplat achieves superior performance in novel view synthesis across most real-world underwater scenes. On the SeaThru-NeRF dataset (levy2023seathru), OceanSplat outperforms the previous state-of-the-art, WaterSplatting (li2024watersplatting), and SeaThru-NeRF-NS (levy2023seathru) by average PSNR margins of 1.05 dB and 2.88 dB, respectively. On the In-the-Wild dataset (tang2024neural), OceanSplat reduces the average LPIPS by $0.021$ compared to 3DGS and surpasses SeaSplat (yang2024seasplat) with average gains of $0.46$ dB in PSNR and $0.014$ in SSIM, along with a $0.024$ reduction in LPIPS.
Figure 4 compares RGB and depth renderings across three underwater scenes. In the Curaçao scene, prior methods suffer from degraded reconstructions due to floating artifacts, whereas our method models the 3D scene with significantly fewer floating artifacts and accurate object representation. In J.G-Redsea scene, prior methods yield plausible RGB images but their depth maps reveal erroneous 3D Gaussians in water regions. In contrast, our method maintains consistent geometry throughout. In the Coral scene, most prior methods (levy2023seathru; kerbl20233d; yang2024seasplat) fail to capture fine-grained coral structures, and WaterSplatting (li2024watersplatting) shows geometry loss and depth holes, while ours reconstructs a faithful 3D scene.
We additionally compare training time, inference FPS, and VRAM usage with prior methods. As shown in Table 3, our method trains faster and consumes less memory than SeaThru-NeRF (levy2023seathru), while remaining competitive among 3DGS-based approaches. Compared to WaterSplatting (li2024watersplatting), our method requires more rasterizations per iteration and least-squares solving for the synthetic epipolar depth prior, resulting in longer training time per-iteration. Nevertheless, it trains over $\times 2$ faster than SeaSplat (yang2024seasplat) while retaining comparable inference speed to other 3DGS-based methods.
@@
We compare the scene reconstruction and restoration performance of SeaThru-NeRF-NS (levy2023seathru), SeaSplat (yang2024seasplat), WaterSplatting (li2024watersplatting), and our proposed OceanSplat using a simulated dataset. As shown in Table 2, for the novel view synthesis task, OceanSplat achieves an average PSNR improvement of 7.39 dB over SeaSplat (yang2024seasplat) and 0.67 dB over WaterSplatting (li2024watersplatting). For scene restoration, our method also demonstrates superior performance, yielding average SSIM gains of 0.048 over SeaSplat (yang2024seasplat) and 0.021 over WaterSplatting (li2024watersplatting). Figure 5 shows that, in both simulated scenes, previous methods assign 3D Gaussians to the scattering medium (e.g., ceilings or walls), resulting in restoration holes, whereas our method accurately represents geometric structures in both underwater and foggy scenes.
@@
Table 5 compares rendering performance and training time across different stereo configurations, including the baseline (li2024watersplatting). Binocular stereo setup $(P_{c}+P_{h})$ improves over baseline, while adding another horizontal stereo view $(P_{c}+P_{h}+P_{h})$ provides minimal gain despite longer training. In contrast, introducing vertical disparity $(P_{c}+P_{h}+P_{v})$ achieves a $0.29$ dB increase in PSNR over the binocular setup. This shows that using horizontally and vertically translated camera poses enhances performance without compromising training time, as shown in Figure 3.
@@ Conclusion
We introduce OceanSplat, a novel framework for enhancing geometric consistency in underwater scene reconstruction. Primarily, we propose trinocular view consistency across translated camera views along two orthogonal axes, effectively aligning 3D Gaussians with scene geometry. In addition, we introduce a synthetic epipolar depth prior derived via triangulation from translated camera views, enabling self-supervised geometric regularization. Lastly, we design a depth-aware alpha adjustment that uses depth and viewing direction to suppress medium-induced floating artifacts. Through these contributions, OceanSplat achieves high-fidelity underwater scene reconstruction. Future work will focus on addressing non-rigid object representations and eliminating the dependency on SfM, thereby broadening the applicability to dynamic underwater scenes.

### v2

@@ Introduction
We evaluate OceanSplat with previous methods, including the official and Nerfstudio versions of SeaThru-NeRF (Levy et al. 2023), as well as 3DGS (Kerbl et al. 2023), SeaSplat (Yang et al. 2024a), and WaterSplatting (Li et al. 2024a). All results were reproduced using the publicly released code. As shown in Table 1, OceanSplat achieves superior performance in novel view synthesis across most real-world underwater scenes. On the SeaThru-NeRF dataset (Levy et al. 2023), OceanSplat outperforms the previous state-of-the-art, WaterSplatting (Li et al. 2024a), and SeaThru-NeRF-NS (Levy et al. 2023) by average PSNR margins of 1.05 dB and 2.88 dB, respectively. On the In-the-Wild dataset (Tang et al. 2024), OceanSplat reduces the average LPIPS by $0.021$ compared to 3DGS and surpasses SeaSplat (Yang et al. 2024a) with average gains of $0.46$ dB in PSNR and $0.014$ in SSIM, along with a $0.024$ reduction in LPIPS.
Figure 4 compares RGB and depth renderings across three underwater scenes. In the Curaçao scene, prior methods suffer from degraded reconstructions due to floating artifacts, whereas our method models the 3D scene with significantly fewer floating artifacts and accurate object representation. In J.G-Redsea scene, prior methods yield plausible RGB images but their depth maps reveal erroneous 3D Gaussians in water regions. In contrast, our method maintains consistent geometry throughout. In the Coral scene, most prior methods (Levy et al. 2023; Kerbl et al. 2023; Yang et al. 2024a) fail to capture fine-grained coral structures, and WaterSplatting (Li et al. 2024a) shows geometry loss and depth holes, while ours reconstructs a faithful 3D scene.
We additionally compare training time, inference FPS, and VRAM usage with prior methods. As shown in Table 3, our method trains faster and consumes less memory than SeaThru-NeRF (Levy et al. 2023), while remaining competitive among 3DGS-based approaches. Compared to WaterSplatting (Li et al. 2024a), our method requires more rasterizations per iteration and least-squares solving for the synthetic epipolar depth prior, resulting in longer training time per-iteration. Nevertheless, it trains over $\times 2$ faster than SeaSplat (Yang et al. 2024a) while retaining comparable inference speed to other 3DGS-based methods.
@@
We compare the scene reconstruction and restoration performance of SeaThru-NeRF-NS (Levy et al. 2023), SeaSplat (Yang et al. 2024a), WaterSplatting (Li et al. 2024a), and our proposed OceanSplat using a simulated dataset. As shown in Table 2, for the novel view synthesis task, OceanSplat achieves an average PSNR improvement of 7.39 dB over SeaSplat (Yang et al. 2024a) and 0.67 dB over WaterSplatting (Li et al. 2024a). For scene restoration, our method also demonstrates superior performance, yielding average SSIM gains of 0.048 over SeaSplat (Yang et al. 2024a) and 0.021 over WaterSplatting (Li et al. 2024a). Figure 5 shows that, in both simulated scenes, previous methods assign 3D Gaussians to the scattering medium (e.g., ceilings or walls), resulting in restoration holes, whereas our method accurately represents scene structures in both underwater and foggy scenes.
@@
Table 5 compares rendering performance and training time across different stereo configurations, including the baseline (Li et al. 2024a). Binocular stereo setup $(P_{c}+P_{h})$ improves over baseline, while adding another horizontal stereo view $(P_{c}+P_{h}+P_{h^{\prime}})$ provides minimal gain despite longer training. In contrast, introducing vertical disparity $(P_{c}+P_{h}+P_{v})$ achieves a $0.29$ dB increase in PSNR over the binocular setup. This shows that using horizontally and vertically translated camera poses enhances performance without compromising training time, as shown in Figure 3.
@@ Conclusion
We introduce OceanSplat, a novel framework for enhancing geometric consistency in underwater scene reconstruction. Primarily, we impose trinocular view consistency across horizontally and vertically translated virtual viewpoints, effectively aligning 3D Gaussians with scene structure. In addition, we introduce synthetic epipolar depth priors derived via triangulation of correspondences from the translated viewpoints, enabling self-supervised depth regularization. Lastly, we propose a depth-aware alpha adjustment that uses depth and viewing direction to suppress medium-induced floating artifacts. Through these contributions, OceanSplat achieves high-fidelity underwater scene reconstruction. Future work will focus on addressing non-rigid object representations and eliminating the dependence on SfM, thereby broadening the applicability to dynamic underwater scenes.